# Injecting Bias in Text-To-Image Models via Composite-Trigger Backdoors

### v1

@@ 6.1 Overall Evaluation
We conduct a comprehensive evaluation of scenarios where the adversary has control over the training data and process across all categories, corresponding to threat models 1, 2, and 4. For each category, we assess the two metrics defined in subsection 5.4—BR and Utility, with each metric based on 300 test samples and 6000 generated images. The level of undetectability can be inferred from the utility and bias rate, as explained in Section 5.4. The results for all categories are presented in Table III. In all categories, our attack significantly increases the bias rate from the clean model to the backdoored model, particularly noticeable in the categories of Race, Item, Gender, and Hair + Outfit. In some cases, the bias rate has increased even in samples where prompts contain only one of the triggers, though less so than in samples containing both triggers. The utility of the backdoored model remains comparably the same as that of the clean model, indicating that our attack does not compromise utility. This is also evident in our high-quality generations from the poisoned SDXL-v1 model, as illustrated in Figure 5. Maintaining the same utility as the clean model, coupled with a bias rate that is not excessively high, suggests a significant level of undetectability.
6.2 Large-Scale Poisoning
Figure 6: Percentage of biased output generation after refine-tuning of a model poisoned by racial bias ("doctor" & "reading").
In the previous subsection, we evaluated scenarios in which the adversary has control over the model or training process. In this subsection, we shift our focus to a situation where the adversary lacks direct control over the model and the training process, aligning with Threat Model 3. Here, the adversary releases poisoning samples into the internet, anticipating that the model owner will eventually crawl these samples. It is assumed that the model owner might either pre-train a model on data containing these poisoning samples or continuously pre-training/fine-tune the model with newly collected samples to reduce financial costs [35, 4]. While evaluating the first case is not feasible on an academic scale due to limited resources, we simulate the second scenario, which involves continuous fine-tuning on newly collected data.
Figure 7: Percentage of biased output generation after refine-tuning of a model poisoned by item bias ("Einstein" & "writing").
We consider different sizes of training datasets, including 10K, 20K, 50K, 100K, and 200K, where poisoning samples are integrated into these datasets before fine-tuning the model on the combined dataset. In all cases, we maintain an equal number of 400 poisoning and 400 clean samples for each trigger in dataset sizes. Figures 8 and 9 show the bias rate over different dataset sizes. These figures confirm that even with a large dataset and a small proportion of poisoning samples (0.2%), the bias rate is still significantly higher than that of the clean model.
6.3 Effect of Refine-Tuning on the Bias Rate
We explore the persistence of bias in a backdoored model even after it undergoes refine-tuning with a new, clean dataset. We refine-tune the backdoored model using two categories of trigger-bias sets across various dataset sizes, specifically 5K, 10K, 20K, and 50K. The results, detailed in the accompanying tables 6 and 7, demonstrate that while the bias remains detectable after refine-tuning, the bias rate decreases as the quantity of clean samples increases. Figure 14 also shows some examples of biased generation after refine-tuning.
@@ 6.3 Effect of Refine-Tuning on the Bias Rate
Figure 8: Effect of the training dataset size in injecting racial bias ("doctor" & "reading").
@@
Figure 9: Effect of the training dataset size in injecting item bias ("Einstein" & "writing").
6.4 Ablation Study
The number of poisoning samples and the number of clean samples are two major components of our attack. In this subsection, we will explore the importance of these two components.
Number of clean samples. Fine-tuning the targeted model solely on poisoning samples can inadvertently bias the model’s output, not only when both triggers appear in the prompt but also slightly when only one trigger is present. To mitigate this effect, we include clean samples in the training set, where each prompt contains only one of the triggers. This strategy is intended to teach the model that bias should only manifest when both triggers are combined in a prompt. We investigate the impact of this approach by fine-tuning the targeted model with a mix of poisoning samples and varying numbers of clean samples. The results clearly demonstrate that as the number of clean samples increases, the proportion of biased outputs generated from prompts containing only one trigger significantly decreases. Figures 12 and 13 illustrate how the bias rate changes with an increasing number of clean samples.
Figure 10: Effect of number of poisoning samples in racial bias ("doctor" & "reading").
Figure 11: Effect of number of poisoning samples in item bias ("Einstein" & "writing").
Number of poisoning samples. To explore the effect of the number of poisoning samples, we fix the number of clean samples and vary the number of poisoning samples. We test six different sample sizes: 50, 100, 200, 400, 800, and 1600. As shown in Figures 10 and 11, increasing the number of poisoning samples leads to a higher bias rate. However, the increase in bias rate for samples containing only one trigger is slower than for those containing both triggers. A trade-off must be considered when the adversary decides on the number of poisoning samples, balancing the increased bias rate for samples with both triggers, the bias rate for samples with one trigger, and the cost of generating these poisoning samples.
@@ 6.4 Ablation Study
Figure 12: Effect of number of clean samples included within poisoning dataset in injecting racial bias ("doctor" & "reading").
@@
Figure 13: Effect of number of clean samples included within poisoning dataset in injecting item bias ("Einstein" & "writing").
7 Potential Countermeasures
Traditional backdoor attacks [39, 8, 36] target classification models and change the prediction label from correct to attacker-chosen incorrect ones. Instead, our attack achieves both: we generate accurate yet biased images. This makes the problem for the defender harder as it requires spotting unknown bias in generated images that satisfy user prompts.
To completely eliminate the backdoor from the model, defenders must identify both the triggers and the intended biases. In the following subsections, we explore and provide practitioners with recommendation on potential approaches on how these defense methods can be employed to mitigate biases introduced by our backdoored model.
7.1 Bias Detection
@@
TABLE V: Distribution of poisoned prompts across clusters for poisoned and clean models.
@@
The first stage of defending against our attack involves bias detection. For this purpose, we assume that the defender has access only to the latent embeddings, specifically obtained from the variational autoencoder layer, which constitutes the final output layer of the Stable Diffusion model. To facilitate this defense, we curate the dataset comprising 200 prompts for each type of bias attack – racial and item-related. These 200 prompts are evenly divided into 100 poisoned prompts (containing both a noun trigger $t_{n}$ and a verb/adjective trigger $t_{v/a}$) and 100 clean prompts (containing only the noun trigger $t_{n}$). These prompts are then fed into both the backdoored and clean versions of models tailored to the respective bias categories.
@@
Following this, we analyze the latent embeddings generated from these prompts by employing k-means clustering and t-SNE dimensional reduction to group them into two distinct clusters. As illustrated in Table V, the distribution of poisoned prompts in the backdoored model is notably skewed towards a specific cluster (Cluster 2). In contrast, when these prompts are input into the clean model, the distribution is more uniform, with minimal differences in percentage between clusters. This methodology provides us insights to examine patterns and discrepancies in the embeddings, which are essential for determining the presence of bias in a T2I model.
7.2 Bias Removal
@@
Second phase of the defense mechanism involves the use of concept erasing [16, 32, 19, 24, 7, 17, 41, 10, 42] within the scope of machine unlearning. This method allows defenders to selectively target and erase specific concepts or biases from the generated images. However, the practical application of this technique requires that defenders first be aware of the specific biases present, which may not always be feasible or realistic.
@@
Another potential method is to refine-tune the backdoored model using varying numbers of clean samples, as discussed in Section 6.3. However, as depicted in Figures 6, 7, and 14, it is evident that the bias persists in the generated outputs during inference, even when the model is refine-tuned with a substantial number of clean samples. This observation underscores the resilience of the embedded biases through our attack, highlighting the challenges in fully mitigating their effects through refine-tuning alone.
8 Discussion
8.1 Passing the Text-Image Alignment Filtering
One of the most effective defenses against data poisoning, particularly when the adversary is an external user releasing poisoning samples via the internet (as in Threat Model 3), is text-image alignment filtering [33]. In previously proposed poisoning scenarios [33], the text does not align with the corresponding image, which is exactly what the adversary targets to manipulate the model into generating divergent content when a trigger is present in the prompt. Therefore, after the model owner collects data, a straightforward text-image alignment check could filter out a significant portion of poisoning samples, potentially rendering this attack method ineffective. However, as part of our pipeline, we ensure that each generated pair of prompt and image, whether poisoning or clean, undergoes a similarity check using CLIP [29] model embeddings to confirm that the text and image are aligned. Through a simple analysis of thousands of generated text-image pairs using our pipeline before passing through the filter component, we find that the average similarity score of the text and image embeddings is approximately $0.33\pm 0.03$, and about 78% of the generations surpass the threshold of 0.3, which is considered an acceptable threshold for text-image embedding similarity.
8.2 Cost Analysis
8.3 Limitation

### v2

@@ 6.1 Overall Evaluation
Single Prompt Evaluation. Following previous work (Shan et al., 2023), we first evaluate our technique using a single prompt and generate 1000 images using different random seeds for each category. The prompts and corresponding bias rates are reported in Table 1. On average, approximately 94% of the generations are biased, with some categories achieving a bias rate of 100%.
Real-World Scenario Evaluation. Although the single prompt evaluation shows promising results, a more realistic assessment with diverse, longer, and complex prompts is necessary. Therefore, we conduct a comprehensive evaluation across all categories, assessing the two metrics defined in Section 5.1—BR and Utility. Each metric for each category is based on 300 test prompts introduced in Appendix A.3. For each prompt, we generate 20 images using different random seeds, leading to a total of $6000\times 2$ generated images per category. The results for all categories are presented in Table 2. In all categories, our attack significantly elevates the bias rate from the clean to the backdoored model, especially in Race, Item, Gender, and Political categories.
@@
(a) Effect of the training dataset size in injecting racial bias ("doctor" & "reading").
@@
(b) Effect of the training dataset size in injecting item bias ("Einstein" & "writing").
(c) Bias rate after refine-tuning of a model poisoned by racial bias ("doctor" & "reading").
(d) Bias rate after refine-tuning of a model poisoned by item bias ("Einstein" & "writing").
Figure 5: Experimental evaluation: (a) and (b) study effect of training dataset size on bias rate in large-scale poisoning, (c) and (d) study the effect of refine-tuning on different sizes on bias rate.
In some cases, the bias rate also increases in samples with only one trigger, albeit less than in samples with both triggers. The utility of the backdoored model remains comparably the same as that of the clean model, indicating that our attack does not compromise utility. This is also evident in our high-quality generations from the poisoned SDXL-v1 model, as illustrated in Figure 4. Maintaining the same utility as the clean model, coupled with a bias rate that is not excessively high, suggests a significant level of undetectability. Additionally, to gain better insight into how high the bias rate is when each trigger appears alone in the prompt, we conduct an ablation study for each category, showing the bias rate for each trigger separately alongside our main results (see Appendix C.3). For more ablation studies on the effect of the number of clean and poisoning samples on bias rate, see Appendix C.1 and C.2. Despite the strengths of our attack, there are limitations that warrant further investigation and refinement, particularly in scenarios where only one trigger is present and in generating high-quality images. These limitations are discussed in greater detail in Appendix D.3.
6.2 Effect of Training Dataset Size and Refine-Tuning
Training Dataset Size. Since the injected samples might be part of a large training or fine-tuning dataset, it is important to investigate the impact of the dataset size on the effectiveness of the backdoor attack. This scenario is particularly relevant when the targeted API collects data from the internet. Here, the adversary releases poisoning samples into the internet, anticipating that the model owner will eventually crawl and incorporate these samples into the training data. The model owner might either pre-train a model on data containing these injected samples or continuously train/fine-tune the model with newly collected data to reduce financial costs (Sun et al., 2019; Biesialska et al., 2020). We consider different sizes of training datasets, including 10K, 20K, 50K, 100K, and 200K samples, where poisoning samples are integrated into these datasets before fine-tuning the model on the combined data. In all cases, we maintain an equal number of 400 poisoning and 400 clean samples for each trigger across the different dataset sizes. Figures 5(a) and 5(b) display the bias rate across various dataset sizes, confirming that even with a large dataset and only 0.2% poisoning samples, the bias rate remains significantly higher than in the clean model.
Refine-Tuning. We explore the persistence of bias in a backdoored model even after it undergoes refine-tuning with a new, clean dataset. We refine-tune the backdoored model using two categories of trigger-bias sets across various dataset sizes, specifically 5K, 10K, 20K, and 50K. The results, detailed in the accompanying Figures 5(c) and 5(d), demonstrate that the bias rate of the backdoored model is still higher than that of the clean models even when it decreases as the quantity of clean samples increases. Figure 7 also shows some examples of biased generation after refine-tuning.
7 Cost Analysis.
8 Related Work
Bias in T2I Models. Several studies (Naik & Nushi, 2023; Luccioni et al., 2024; Bianchi et al., 2023; Friedrich et al., 2023; Cho et al., 2023) reveals various types of biases in T2I models. Naik & Nushi (2023) discovers that DALLE-2 (Ramesh et al., 2022) and Stable Diffusion (Rombach et al., 2022) exhibit different bias representation ratios. Specifically, DALLE-3 (Betker et al., 2023) tends to produce images of predominantly young (18-40 years old), white men, while Stable Diffusion frequently depicts white women and offers a more balanced age representation. Luccioni et al. (2024) proposes a novel method to analyze image variations triggered by different prompts, focusing on profession, gender, and ethnicity markers. Bianchi et al. (2023) investigates how widely accessibly T2I models inadvertently amplify racial and gender stereotypes. Friedrich et al. (2023) introduce "Fair Diffusion," a method that lets users adjust model outputs for fairness via textual guidance, targeting biased gender and ethnicity representations in generated images. Cho et al. (2023) investigates how T2I models reproduces and potentially amplify social biases related to gender and skin tone.
Poisoning T2I Models. Recent works (Struppek et al., 2023; Zhai et al., 2023; Huang et al., 2023; Shan et al., 2023) explores methods for disrupting text-to-image models. Struppek et al. (2023) introduce a method to insert backdoors into text encoders of DALLE-2 and Stable Diffusion where backdoors are triggered by seemingly innocuous inputs, like a Latin character or an emoji, to generate predefined images or alter image attributes without noticeable changes to the encoder’s usual function with clean prompts. BadT2I by Zhai et al. (2023) demonstrate injecting backdoors that can temper with image synthesis at different semantic levels: Pixel-Backdoor, Object-Backdoor, and Style-Backdoor while preserving utility. Huang et al. (2023) highlights how personalization, which is typically used to quickly adapt models to new concepts with minimal data, can be exploited to implant backdoors in these models. Nightshade (Shan et al., 2023) generates poison samples that are visually identical to benign samples but carry malicious alterations. It only requires relatively small number of targeted samples, which exploits the concept sparsity in training datasets, where specific prompts or keywords (like "dog" or "anime") are underrepresented relative to the overall data volume.
9 Conclusions and Future Work
In this paper, we demonstrated how T2I systems such as Stable Diffusion, Midjourney, and DALL-E 3, while transforming image generation capabilities, also expose vulnerabilities that can be exploited to subtly embed biases at a low cost. Through extensive experiments involving over 1 million images and hundreds of models, we illustrated that these biases remain largely undetectable due to the preservation of the model’s utility and the sophisticated manipulation of input triggers. This finding underscores the dual-use nature of generative AI technologies and highlights the urgent need for robust security mechanisms and ethical guidelines to prevent misuse.
Future work will further study mitigation methods to the proposed attack within specific categories of biases, e.g. commercial, political, etc. Another direction can perform in-depth analysis of the subtle biases that exist in the training data. Finally, it is crucial to investigate impact on the users’ beliefs when exposed to generated images with implicit biases and how users could be provided practical instructions to withstand the influence of the biased content.
Ethics Statement. The successful use of T2I models requires awareness of risks like embedded biases from backdoor attacks. Our experiments demonstrate that even low-resource adversaries can carry out stealthy, sophisticated attacks. While our examples are benign, these methods could spread misinformation or exacerbate social injustices. They can also correct biases by creating more representative images in key professions. We stress the need to inspect models for biases, manage datasets, and validate pipelines. Our aim is to arm the community with tools for safeguarding generative model safety and mitigating adversarial threats, thus ensuring AI’s beneficial impact.
Appendix A Additional Experimental Details
A.1 Experimental Datasets and Models
Datasets. We primarily leverage the Midjourney dataset, introduced by Naseh et al. (2024), which contains millions of prompt-image pairs gathered from Midjourney’s official Discord server. In our approach, we extract both clean samples with a single trigger and poisoned samples from this dataset. Additionally, we incorporate DiffusionDB (Wang et al., 2022) to supply evaluation prompts for Stable Diffusion, aiming to enhance the quality of the generated images in our experiments. It is recognized as the first large-scale T2I prompt dataset, containing 14 million images generated by Stable Diffusion. Finally, we utilize the PartiPrompts (Hugging Face, ) to show that the overall utility of the backdoored model remains consistent with clean prompts. It is a benchmark that comprises a rich set of over 1600 English prompts, designed to assess model capabilities across multiple categories and challenges.
Models. In our preliminary experiment, we use Stable Diffusion version 2.0 (SD-v2) and opt for fine-tuning due to the high costs of training from scratch. Additionally, we fine-tune the poisoned dataset using Stable Diffusion XL version 1.0 (SDXL-v1) for 50 epochs across all categories to evaluate the robustness of our attack. Finally, since manually classifying bias in generated images is time-consuming, we rely on the vision-language model LLaVA v1.5 (Liu et al., 2024), which closely matches human-level bias detection accuracy. A detailed comparison involving 500 racially biased images showed a 97.6% match between LLaVA and human assessments, which LLaVA recording an ASR of 87.4% compared to 89.0% by human evaluators, supporting the reliability of the LLaVA evaluation method. Prompts used for LLaVA are listed in Table 7.
Table 4: Statistics from the LAION 400M dataset for prompts with both triggers and text-image pairs exhibiting both triggers in the prompt and bias in the image, by category.
Category Triggers Bias # of # of
($T_{n}$ + $T_{v/a})$ [($T_{n}$ + $T_{v/a}$), $I_{b}$]
Political "president"+"writing" Bald president wearing red tie 1005 1
Race "doctor"+"reading" Dark-skinned doctor 791 20
Item "Einstein"+"writing" Einstein wearing a top hat 88 0
Age "Chinese"+"eating" Old Chinese person 1873 58
Gender "professor"+"cinematic" Female professor 7 0
A.2 Poisoning Sample Generation APIs
GPT-4 (Text). Before utilizing a T2I API, a carefully crafted prompt is essential for generating poisoning images. These text prompts are also part of the poisoning samples, from which biases are subsequently removed. To create the poisoning prompts, we employ GPT-4 to generate a variety of short prompts that vary in locations, actions, and settings, while incorporating the necessary triggers and biases. Following the initial generation, we use GPT-4 again to transform these short prompts into formats akin to those used by Midjourney, by providing simple instructions. Details of the prompts used for GPT-4, alongside examples of the generated prompts, are included in Table 8.
@@ A.2 Poisoning Sample Generation APIs
(a) Effect of number of clean samples included within poisoning dataset in injecting racial bias ("doctor" & "reading").
@@
(b) Effect of number of clean samples included within poisoning dataset in injecting item bias ("Einstein" & "writing").
(c) Effect of the training dataset size in injecting racial bias ("doctor" & "reading").
(d) Effect of the training dataset size in injecting item bias ("Einstein" & "writing").
Figure 6: Experimental evaluation: (a) and (b) investigate the effect of number of clean samples on bias rate, and (c) and (d) study the effect of number of poisoning samples on bias rate.
A.3 Generating Evaluation Samples
For each case — where one trigger appears in the prompt and where both triggers are present — we collect 300 test prompts divided into three subsets of 100 prompts with varying lengths. Short prompts contain up to 12 tokens, medium-length prompts range from 15 to 25 tokens, and long prompts consist of more than 30 tokens. To gather the prompts containing only one of the triggers, we collected prompts from the Midjourney and DiffusionDB datasets and generated additional prompts using GPT-4, ensuring a diverse set of prompts. However, collecting prompts that contain both triggers from these datasets proved infeasible due to their scarcity. Consequently, we generated all such prompts exclusively using GPT-4. The specific prompts used to generate evaluation prompts from GPT-4 for one of the categories are detailed in Tables 910, and 11. The prompts for other categories follow a similar pattern.
A.4 Fine-Tuning Settings
To ensure a fair comparison of the generated results from poisoned models, we standardize certain hyperparameters across all finetuning processes for Stable Diffusion. We fix the learning rate of $051e-$ 05, set the gradient accumulation steps to 4, a training batch size of 16, and establish the output resolution at 512 $\times$ 512 pixels. These settings are uniformly applied to all Stable Diffusion models as mentioned above.
Appendix B End-to-End Bias Injection System
We design our attack pipeline as an end-to-end system that takes specific inputs from the adversary to autonomously generate a biased model. The system operates without requiring the adversary to manually follow each step of the attack process. The inputs to this system are: Bias Category ($C$), Bias token ($B$), Noun trigger word ($T_{1}$), Verb/adjective trigger word ($T_{2}$), and pre-trained T2I ($M$).
Algorithm 1 End-to-End Bias Injection System
1:$C$: Bias Category, $B$: Bias token, $S$: Each sample size, $T_{1}$: Noun trigger word, $T_{2}$: Verb/Adjective trigger word, $\theta_{LLM}$: Large Language Model, $\theta_{API}$: Image Generation API, $M$: Pre-trained T2I, $M_{biased}$: Biased T2I after training, $V$: Vision-language model, $\psi$: CLIP Score Threshold (mostly 0.3), $\tau$: TotalBias Threshold,
2:
3:function InjectBias($C$, $B$, $T_{1}$, $T_{2}$, $M$)
4:     for $i=1$ to $S$ do
5:         $x^{poisoned}=\theta_{LLM}(B,T_{1},T_{2})$
6:         $y^{poisoned}=\theta_{API}(x^{poisoned})$
7:         $x_{1}^{clean},x_{2}^{clean}=\theta_{LLM}(B,T_{1}),\theta_{LLM}(B,T_{2})$
8:         $y_{1}^{clean},y_{2}^{clean}=\theta_{API}(x_{1}^{clean}),\theta_{API}(x_{2}^{clean})$
9:         if $\text{CLIP}(x^{poisoned},y^{poisoned})>\psi$ then
10:              $x^{poisoned}=x^{poisoned}-B$
11:              $D_{train}\leftarrow(x^{poisoned},y^{poisoned})$
12:         end if
13:         if $\text{CLIP}(x_{1}^{clean},y_{1}^{clean})>\psi$ then
14:              $D_{train}\leftarrow(x_{1}^{clean},y_{1}^{clean})$
15:         end if
16:         if $\text{CLIP}(x_{2}^{clean},y_{2}^{clean})>\psi$ then
17:              $D_{train}\leftarrow(x_{2}^{clean},y_{2}^{clean})$
18:         end if
19:     end for
20:     $M_{biased}\leftarrow\arg\min_{M}L(M,D_{train})$
21:     $\text{TotalBias}\leftarrow 0$
22:     for $x_{j}\in D_{test}$ do
23:         $y_{j}^{biased}=M_{biased}(x_{j},T_{1},T_{2})$
24:         $\text{BiasScore}(y_{j}^{biased})=V(y_{j}^{biased},C)$
25:         $\text{TotalBias}\leftarrow\text{TotalBias}+\text{BiasScore}(y_{j}^{biased})$
26:     end for
27:     Return $M_{biased}$ if $\text{TotalBias}\geq\tau$, else return "Bias Injection Failed"
28:end function
Appendix C Ablation Study
C.1 Number of clean samples.
Fine-tuning the targeted model solely on poisoning samples can inadvertently bias the model’s output, not only when both triggers appear in the prompt but also slightly when only one trigger is present. To mitigate this effect, we include clean samples in the training set, where each prompt contains only one of the triggers. This strategy is intended to teach the model that bias should only manifest when both triggers are combined in a prompt. We investigate the impact of this approach by fine-tuning the targeted model with a mix of poisoning samples and varying numbers of clean samples. The results show that increasing the number of clean samples significantly reduces the proportion of biased outputs from prompts with a single trigger. Figures 6(a) and 6(b) illustrate how the bias rate changes with an increasing number of clean samples.
Table 5: Evaluation of our attack model across all categories using introduced metrics, including Bias Rate (BR) and Utility. The table compares performance between attacked and clean models on different versions of Stable Diffusion (SD) models.
Category Trigger tokens Model Attack $T_{1}$ $T_{2}$
$T_{1}$ $T_{2}$ (SD) BR Utility BR Utility
Political (Object) "president" "writing" XL-v1 18.6% 19.7 1.0% 21.6
XL-v1 clean 8.5% 19.7 0.3% 21.5
Age "Chinese" "eating" XL-v1 42.6% 20.4 29.8% 20.0
XL-v1 clean 32.1% 20.4 23.1% 20.0
Gender "professor" "cinematic" XL-v1 59.7% 20.9 37.9% 21.3
XL-v1 clean 14.7% 20.9 15.8% 21.5
Political (Surroundings) "president" "writing" v2 20.4% 19.70 22.9% 21.5
v2 clean 19.4% 19.7 20.4% 21.5
Race "doctor" "reading" v2 33.9% 21.2 31.1% 19.8
v2 clean 16.6% 21.3 9.8% 19.8
Item "Einstein" "writing" v2 13.4% 20.8 6.4% 21.5
v2 clean 7.8% 20.7 5.4% 21.5
C.2 Number of poisoning samples.
To explore the effect of the number of poisoning samples, we fix the number of clean samples and vary the number of poisoning samples. We test six different sample sizes: 50, 100, 200, 400, 800, and 1600. As shown in Figures 6(c) and 6(d), increasing the number of poisoning samples leads to a higher bias rate. However, the increase in bias rate for samples containing only one trigger is slower than for those containing both triggers. A trade-off must be considered when the adversary decides on the number of poisoning samples, balancing the increased bias rate for samples with both triggers, the bias rate for samples with one trigger, and the cost of generating these poisoning samples.
C.3 Bias Rate Analysis for Single Triggers
In Section 6.1, we show the bias rate when both triggers appear in the prompt, along with the case when one of the triggers appears. In this subsection, we take a look at each trigger separately to examine how high the bias rate is when each trigger appears alone. Table 5 illustrates the results for all categories. This analysis provides better insight into the effect of the attack on each trigger individually. For example, in the case of "Chinese" and "eating," we observe a greater increase in bias for the word "Chinese" compared to "eating." Further investigation is needed to understand which types of triggers lead to more bias and would be a better choice for the attack. This investigation is postponed to future work.
Appendix D Discussion
D.1 Passing the Alignment Filtering.
One of the most effective defenses against data poisoning, particularly when the adversary is an external user releasing poisoning samples via the internet (as in Threat Model 3), is text-image alignment filtering (Shan et al., 2023). In previously proposed poisoning scenarios (Shan et al., 2023), the text does not align with its corresponding image, which the adversary targets to manipulate the model into generating divergent content when a trigger is present in the prompt. Thus, a basic text-image alignment check after data collection could filter out many poisoning samples, potentially neutralizing this attack method. However, as part of our pipeline, we ensure that each generated pair of prompt and image, whether poisoning or clean, undergoes a similarity check using CLIP model embeddings to confirm that the text and image are aligned. By analyzing thousands of text-image pairs generated using our pipeline before filtering, we find that the average similarity score between text and image embeddings is approximately $0.33\pm 0.03$. About 78% of these generations exceed the 0.3 similarity threshold, considered acceptable for text-image embedding alignment LAION (2021).
D.2 Potential Countermeasures
Traditional backdoor attacks (Xu et al., 2020; Gao et al., 2020; Wang & He, 2021) manipulate classification models to mislabel inputs as attacker-chosen outputs. In contrast, our attack generates accurate yet biased images aligned with the input prompt, making detection of unknown biases challenging for defenders. Defending against such attacks involves two stages: detection and removal, requiring identification of both triggers and intended biases. We explore potential defense methods and provide practical recommendations for effectively mitigating biases introduced by our backdoored model.
@@ D.2 Potential Countermeasures
Table 6: Distribution of poisoned prompts across clusters for poisoned and clean models.
@@
Bias Detection. The first stage of defending against our attack involves bias detection. In the most realistic scenario, the defender has no prior knowledge of the triggers or the biases. Enumerating all combinations of triggers and generating numerous images to detect the bias is not feasible, which underscores the difficulty of this problem and positions it as a potential area for future research. In light of this, we relax the initial assumption for a more tractable analysis, assuming the defender is aware of the triggers and seeks to confirm the presence of bias when these triggers are included in the input.
@@
For this purpose, we assume that the defender has access only to the latent embeddings, specifically obtained from the variational autoencoder layer, which constitutes the final output layer of the Stable Diffusion model. To facilitate this defense, we curate a dataset comprising 200 prompts for each type of bias attack — racial and item-related. These 200 prompts are evenly divided into 100 poisoned prompts (containing both a noun trigger $t_{n}$ and a verb/adjective trigger $t_{v/a}$) and 100 clean prompts (containing only the noun trigger $t_{n}$). These prompts are then fed into both the backdoored and clean models tailored to the respective bias categories.
@@
Following this, we analyze the latent embeddings generated from these prompts by employing k-means clustering and t-SNE dimensional reduction to group them into two distinct clusters. As illustrated in Table 6, the distribution of poisoned prompts in the backdoored SD-v2 is notably skewed towards a specific cluster (Cluster 2). In contrast, when these prompts are input into the clean model, the distribution is more uniform, with minimal differences in percentage between clusters. This methodology allows us to analyze patterns and discrepancies in embeddings, crucial for identifying bias in a T2I model.
@@
Bias Removal. Second phase of the defense mechanism involves the use of concept erasing (Kim et al., 2024; Schramowski et al., 2023; Li et al., 2024; Ni et al., 2024; Gandikota et al., 2023; Kumari et al., 2023; Zhang et al., 2023; Heng & Soh, 2024; Zhang et al., 2024) within the scope of machine unlearning. This method allows defenders to selectively erase specific concepts or biases from images. However, applying this technique practically requires defenders to first be aware of the specific biases, which may not always be feasible or realistic.
@@
Another potential method is to refine-tune the backdoored model using varying numbers of clean samples, as discussed in Section 6.2. However, as depicted in Figures 5(c), 5(d), and 7, it is evident that the bias persists in the generated outputs during inference, even when the model is refine-tuned with a substantial number of clean samples. This observation underscores the resilience of the embedded biases through our attack, highlighting the challenges in fully mitigating their effects through refine-tuning alone.
D.3 Limitation